A turtle accelerates at a rate of 2 m/sfor 15 s, starting from no motion. What is the turtle's final velocity?




1
Expert's answer
2022-04-21T13:01:02-0400

Given:

"v_i=0\:\rm m/s"

"a=2\:\rm m/s^2"

"t=15\:\rm s"


The final velocity

"v_f=v_i+at\\
=0+2*15=30\:\rm m/s"